Solve for the remaining 5 trig functions; tanθ=-1/4, θ is in quadrant IV
Answer by stanbon(75887)

Solve for the remaining 5 trig functions; tanθ=-1/4, θ is in quadrant IV
-----------
In QIV x is positive and y is negative.
-----------
Since tan = y/x, y = -1 and x = 4
Then r = sqrt[1^2 + 4^2] = sqrt(17)
----------------------------
sin = y/r = -1/sqrt(17)
cos = x/r = 4/sqrt(17)
tan = y/x = -1/4
-----
csc = r/y = -sqrt(17)
sec = r/x = sqrt(17)/4
cot = x/y = -4
---------------------------
Cheers,
Stan H.
-------------------